# Device-Independent Certification of Multipartite Distillable Entanglement

### v1

@@ V.2 EAT Channels
Consider Condition 1. We can think of $R_{j}$ as a source distributing the states across the network. As far we are concerned, it is an arbitrary quantum system. $W_{j}$ is the classical register associated with determining whether the MABK game has been won. $W_{j}$ is finite-dimensional as it can take only three values. $S_{j}$ consists of all the systems involved in the protocol, except for the $\hat{A}_{[M^{\prime}],\,j}$ for some $M^{\prime}\in[M-1]$.
For Condition 2, we can see that $W_{j}$ is determined using only the classical values $A_{i,j}$ and $X_{i,j}$ and does not affect the classical and quantum registers.
Finally, we are left with Conditions 3. Checking these conditions is more involved. First, we note that Protocol 1 makes use of the MABK inequality, which involves two inputs and two outputs for all parties involved. This allows us to apply Jordan’s lemma on every party. Now, we recall Jordan’s lemma [44]:
@@
Let $\hat{O}_{0}$ and $\hat{O}_{1}$ be two Hermitian operators with eigenvalues $-1$ and $+1$. There exists a basis in which both operators are block-diagonal, in blocks of dimension $2\times 2$ at most.
@@
If we act with the projection $\Pi_{d_{i,j}}$ on $\rho_{\hat{A}_{[M],j}}$, then we obtain the index $d_{i,j}$ of the corresponding Jordan block. Hence, after the application of the projection, the system possessed by each party is a two-dimensional system. This is true of $\hat{A}_{i,j}$ especially, which satisfies Condition 3. The post-measurement state can be written as follows:
@@
Using these projections, we produce Protocol 2. Protocol 2 is a device-dependent version of Protocol 1 but has the same winning probability as Protocol 1. This is due the fact that, regardless of the value of $T_{j}$ and for any $d_{j}=d_{1,j}\cdots d_{M,j}$, the underlying state can thought of as being in the following state [45, Theorem 1]:
@@
where $\mathbf{i}^{2}=-1$, $s^{d_{j}}_{u},\lambda^{d_{j}}_{0,u}\in[0,1]$, and $\psi_{0,u},\psi_{1,u}$ are defined in Definition 1. Hence, the projective measurement detailed earlier will neither change the winning probability nor the amount of multipartite entanglement produced by the source. (This is true for any two-input two-output, or ($M$, 2, 2), full-correlator Bell inequality [45, Theorem 1], like the MABK inequality we consider here.)
@@ V.3 Max-Tradeoff Function
To calculate this upper bound, we will use the fact that the conditional entropy is concave [46]:
@@
where $\{p(x)\}_{x}$ is a probability distribution, $\{\rho^{x}_{AB}\}_{x}$ is a set of states, and $\bar{\rho}_{AB}\coloneqq\sum_{x}p(x)\rho_{AB}^{x}$. We will also make use of the bipartition between systems represented by the vertical bar in (39), which means that we can make use of local operations with respect to this bipartition. Given a particular $D_{[M],j}$ at the end of the $j$-th round of Protocol 2, the state $\tilde{\rho}^{d_{j}}$ in (38) has a simple structure in the GHZ basis. We can rewrite the GHZ basis, from Definition 1, in the following fashion:
@@
Within the partition denoted by $u$, we can choose one qubit to be the control qubit, $u_{1}$, and then apply CNOTs on all the other qubits. We apply similar set of operations on the partition denoted by $w$ with the first qubit containing $w_{1}$ as the control qubit. These operations will transform the $M$-partite GHZ basis states into the following state:
@@
These lemmas can be applied to Protocol 2 in the following manner. Note that the following equality holds:
$\sum_{x}p(x)H(A|B)_{\rho^{x}}=H(A|BX)_{\tilde{\rho}},$ (47)
where $\{p(x)\}_{x}$ is a probability distribution, and $\{\rho^{x}_{AB}\}_{x}$ is a set of states, $\tilde{\rho}_{ABX}\coloneqq\sum_{x}p(x)\rho_{AB}^{x}\otimes|x\rangle\!\langle x|$, and $\{|x\rangle\}_{x}$ is an orthonormal basis. Using this, we find that
$\sup_{\sigma\in\Sigma_{p}}\Bigg{[}\mathrm{Pr}(T_{j}=0)H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}(\sigma)}$
$+\mathrm{Pr}(T_{j}=1)H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}(\sigma)}\Bigg{]}.$ (48)
Note that $T_{j}=1$ corresponds to a round where we measure that system to play the MABK game and each $\hat{A}_{i,j}$ is in a deterministic state for all $i\in[M]$. So,
$\sup_{\sigma\in\Sigma_{p}}H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\,T_{j}\right)_{\mathcal{N}_{j}(\sigma)}=\\ (1-\gamma)\sup_{\sigma\in\Sigma_{p}}H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}(\sigma)}.$ (49)
Using Lemma 1 and Lemma 2, we find that
@@
Proof. When $M$ is even, the MABK game winning probability and the MABK violation in (13) are related as follows: $\beta_{M}=8\omega-4$. Substituting $\beta_{M}=8\omega-4$ in (50), we get the required result.
@@
Proof. When $M$ is odd, the MABK game winning probability and the MABK violation in (13) are related as follows: $\beta_{M}=8\sqrt{2}\omega-2\sqrt{2}$. Substituting $\beta_{M}=8\sqrt{2}\omega-2\sqrt{2}$ in (50), we get the required result.
Figure 1: Plot of $g^{e}(\omega)$, defined in (52), for $\omega\in[p^{e}_{\min},p^{e}_{\max}]$ defined in (22). The solid line corresponds to $g^{e}(\omega)$. The dashed line corresponds to $g^{e}(\omega)=0$. Note that $g^{e}(\omega)$, defined in (54), for $\omega\in[p^{o}_{\min},p^{o}_{\max}]$, defined in (23), looks very similar.
We plot $g_{e}(\omega)$ in Figure 1. These theorems serve as the basis for our max-tradeoff function. Using Theorems 6 and 7 to define a max-tradeoff function for all $p$ with $p(1)/\gamma\in[\,p_{\min},p_{\max}]$. Define a function $f$ in the following fashion: when $M$ is even
@@ V.4 Applying the EAT
We can finally apply the EAT, stated in Theorem 4, to derive an upper bound on the smooth max-entropy. By obtaining the exact form of $f_{\max}$, we have the calculated $t$ on the right-hand side of (32). We need to calculate (33). We know that the dimension of all quantum registers involved is two, which means that $d_{O_{i}}=2M^{\prime}$. From (57), we get $\left\|\triangledown f_{\max}\right\|_{\infty}\leq a(p_{t})$. According to Condition 2 in Definition 3, for soundness, we require that, for every source $\phi$ and measurement device, Protocol 1 either certifies a minimum amount of GHZ entanglement or aborts with probability greater than $1-\varepsilon_{\operatorname{snd}}$ when applied on $\sigma$. So, the probability of Protocol 1 not aborting is negligible i.e., $\Pr[\Omega]_{\tau}\leq\varepsilon_{\operatorname{snd}}$. Also, note that $\varepsilon_{\operatorname{smo}}$ is chosen such that the fidelity constraint in (2) is satisfied. When Protocol 2 does not abort, we can define the following quantity, which corresponds to the right-hand side of (32):
$\eta(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},p_{t},\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},M^{\prime},M)=\\ n\cdot f_{\max}\!\left(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},p_{t},M\right)+2\left(\log_{2}(1+2M^{\prime})\sqrt{n}+\lceil a(p_{t})\rceil\right)\\ \times\sqrt{1-2\log_{2}\left(\varepsilon_{\mathrm{smo}}\cdot\varepsilon_{\operatorname{snd}}\right)},$ (60)
@@
$\eta_{\mathrm{opt}}\left(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},M^{\prime},M\right)=\\ \min_{p_{t}:p_{\min}<\frac{p_{t}(1)}{\gamma}<p_{\max}}\eta\left(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},p_{t},\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},M^{\prime},M\right).$ (61)
@@
For every source and all measurement devices in the setting detailed earlier, let $\rho$ be the state generated using Protocol 2, $\Omega$ the event that Protocol 2 does not abort, and $\rho_{|\Omega}$ the state conditioned on $\Omega$. Then, for all $\varepsilon_{\operatorname{snd}},\varepsilon_{\operatorname{smo}}\in(0,1)$, either Protocol 2 aborts with probability greater than $1-\varepsilon_{\operatorname{snd}}$ or
@@
Now we have an upper bound on the conditional smooth max-entropy. We will use this bound to show that Protocol 1 is sound. Note that all the results discussed so far are applicable in general to every $M$-partite scenario.
@@ V.5 Soundness
In this subsection, we are concerned with the quantity in (30). We need to bound this quantity in order to certify the multipartite entanglement distillation rate from a source in a device-independent manner. We use the preceding theorems to certify the multipartite entanglement distillation rate from a source. This brings us to our final theorem, which is as follows.
@@
For all $\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},\varepsilon_{\mathrm{cmp}}\in[0,1]$, Protocol 1 is a DIMEC protocol with
@@
    $E_{D}^{n,\varepsilon_{\operatorname{snd}}}(\rho_{|\Omega})\geq\\ \frac{-\eta_{\mathrm{opt}}\left(\omega_{\mathrm{exp}},\varepsilon_{\operatorname{snd}},\varepsilon_{\mathrm{smo}},M-1,M\right)+2\log\varepsilon_{\operatorname{smo}}}{M-1},$ (63)
@@
$E_{D}^{n,\varepsilon_{\operatorname{snd}}}(\rho_{|\Omega})\\ \geq\frac{-\eta_{\mathrm{opt}}\left(\omega_{\mathrm{exp}},\varepsilon_{\operatorname{snd}},\varepsilon_{\mathrm{smo}},M-1,M\right)}{M-1}+\frac{2\log\varepsilon_{\operatorname{smo}}}{M-1},$ (67)
@@ VI Conclusion
In this work, we define DI $M$-partite entanglement distillation certification protocol consisting of completeness and soundness conditions. We then propose a protocol and show that the proposed protocol is indeed a DI multipartite entanglement distillation certification protocol. Specifically, we prove that Protocol 1 is indeed an DI multipartite entanglement distillation certification protocol. Our proposed protocol is centered around the MABK inequality [33, 34, 35] which is critical to proving the completeness and soundness conditions. To prove the soundness condition, we use the entropy accumulation theorem [36], the structure of the MABK inequality [33, 34, 35] and the multipartite entanglement distillation protocols and rates described in [40, 41].
For future works, we are interested in using other Bell-type inequalities that involve more two binary measurements or measurement that have more than two outcomes. In this work, we have focused entirely on GHZ states and leaving out the W state. The W and GHZ state are not inter-convertible using LOCC alone [48]. Hence, our analysis here for GHZ states does not immediately hold for W states, but it is an interesting extension. It is also worth noting that there are infinitely many classes of genuinely multipartite entangled states that involve four or more parties [49]. It will be interesting to see what a unified certification protocol for such classes of states might look like.

### v2

@@ V.2 EAT Channels
Consider Condition 1 of Definition 8. We can think of $R_{j}$ as a source distributing the states across the network. As far we are concerned, it is an arbitrary quantum system. $W_{j}$ is the classical register associated with determining whether the MABK game has been won. $W_{j}$ is finite dimensional as it can take only three values. $S_{j}$ consists of all the systems involved in the protocol, except for the system $\hat{A}_{[M^{\prime}],\,j}$ for some $M^{\prime}\in[M-1]$.
For Condition 2 of Definition 8, we can see that $W_{j}$ is determined using only the classical values $A_{i,j}$ and $X_{i,j}$ and does not affect the classical and quantum registers.
Finally, we are left with Condition 3 of Definition 8. Checking these conditions is more involved. First, we note that Protocol 1 makes use of the MABK inequality, which involves two inputs and two outputs for all parties involved. This allows us to apply Jordan’s lemma on every party. Now, we recall Jordan’s lemma [45]:
@@
Let $\hat{O}_{0}$ and $\hat{O}_{1}$ be two Hermitian operators with eigenvalues $-1$ and $+1$. Then there exists a basis in which both operators are block diagonal, in blocks of dimension $2\times 2$ at most.
@@
Let $\Pi_{d_{i,j}}$ denote the projection onto the $d_{i,j}$th block. If we act with the projection $\Pi_{d_{i,j}}$ on $\rho_{\hat{A}_{[M],j}}$, then we obtain the index $d_{i,j}$ of the corresponding Jordan block. Hence, after the application of the projection, the system possessed by each party is a two-dimensional system. This is true of $\hat{A}_{i,j}$ especially, which satisfies Condition 3. After party $i$ performs the measurement $\{\Pi_{d_{i,j}}\}_{d_{i,j}}$ for round $j$, for all $i\in[M]$, the post-measurement state is as follows:
@@
Using these projections, we produce Protocol 2. Protocol 2 is a device-dependent version of Protocol 1 but has the same winning probability as Protocol 1. This is due the fact that, regardless of the value of $T_{j}$ and for any $d_{j}=d_{1,j}\cdots d_{M,j}$, the underlying state can thought of as being in the following state [46, Theorem 1]:
@@
where $\mathbf{i}^{2}=-1$, $s^{d_{j}}_{u},\lambda^{d_{j}}_{0,u}\in[0,1]$, and $\psi_{0,u},\psi_{1,u}$ are defined in Definition 1. Hence, the projective measurement detailed earlier will neither change the winning probability nor the amount of multipartite entanglement produced by the source. (This is true for any two-input two-output, or ($M$, 2, 2), full-correlator Bell inequality [46, Theorem 1], like the MABK inequality we consider here.)
@@ V.3 Max-Tradeoff Function
We will first simplify (39). Note that the following equality holds:
$\sum_{x}p(x)H(A|B)_{\rho^{x}}=H(A|BX)_{\tilde{\rho}},$ (41)
where $\{p(x)\}_{x}$ is a probability distribution, $\{\rho^{x}_{AB}\}_{x}$ is a set of states, $\tilde{\rho}_{ABX}\coloneqq\sum_{x}p(x)\rho_{AB}^{x}\otimes|x\rangle\!\langle x|$, and $\{|x\rangle\}_{x}$ is an orthonormal basis. Using this, we find that
$\sup_{\sigma\in\Sigma_{p}}\Bigg{[}\mathrm{Pr}(T_{j}=0)H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}^{0}(\sigma)}$
$+\mathrm{Pr}(T_{j}=1)H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}^{1}(\sigma)}\Bigg{]},$ (42)
where $\mathcal{N}_{j}^{0}(\sigma)$ corresponds to the state when $T_{j}=0$, and $\mathcal{N}_{j}^{1}(\sigma)$ corresponds to the state when $T_{j}=1$. Note that $T_{j}=1$ corresponds to a round where we apply measurements to play the MABK game and each $\hat{A}_{i,j}$ is in a deterministic state for all $i\in[M]$. This implies that
$H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}^{1}(\sigma)}=0.$ (43)
Hence,
$\sup_{\sigma\in\Sigma_{p}}H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\,T_{j}\right)_{\mathcal{N}_{j}(\sigma)}=\\ (1-\gamma)\sup_{\sigma\in\Sigma_{p}}H\!\left(\hat{A}_{[M^{\prime}],j}\middle|\hat{A}_{[M^{\prime}+1,M],j}\,(ADX)_{[M],j}\right)_{\mathcal{N}_{j}^{0}(\sigma)}.$ (44)
To calculate an upper bound on the above quantity, we will use the fact that the conditional entropy is concave [47]:
@@
where the notation is the same as in (41) and $\bar{\rho}_{AB}\coloneqq\sum_{x}p(x)\rho_{AB}^{x}$. We will also make use of the bipartition between systems represented by the vertical bar in (39), which means that we can make use of local operations with respect to this bipartition. Given a particular $D_{[M],j}$ at the end of the $j$th round of Protocol 2, the state $\tilde{\rho}^{d_{j}}$ in (38) has a simple structure in the GHZ basis. We can rewrite the GHZ basis, from Definition 1, in the following fashion:
@@
Within the partition denoted by $u$, we can choose one qubit to be the control qubit, $u_{1}$, and then apply CNOTs on all the other qubits. We apply similar set of operations on the partition denoted by $w$ with the first qubit containing $w_{1}$ as the control qubit. These operations transform the $M$-partite GHZ basis states into the following state:
@@
These lemmas can be applied to Protocol 2 in the following manner. Using Lemma 1 and Lemma 2, we find that
@@
Proof. When $M$ is even, the MABK game winning probability and the MABK violation in (12) are related as follows [37]: $\beta_{M}=8\omega-4$. Substituting $\beta_{M}=8\omega-4$ in (51), we get the required result.
@@
Proof. When $M$ is odd, the MABK game winning probability and the MABK violation in (12) are related as follows [37]: $\beta_{M}=8\sqrt{2}\omega-2\sqrt{2}$. Substituting $\beta_{M}=8\sqrt{2}\omega-2\sqrt{2}$ in (51), we get the required result.
These theorems serve as the basis for our max-tradeoff function. Using Theorems 6 and 7 to define a max-tradeoff function for all $p$ with $p(1)/\gamma\in[\,p_{\min},p_{\max}]$. Define a function $f$ in the following fashion: when $M$ is even
@@ V.4 Applying the EAT
We can finally apply the EAT, stated in Theorem 4, to derive an upper bound on the conditional smooth max-entropy. By obtaining the exact form of $f_{\max}$, we have the calculated $t$ on the right-hand side of (32). We need to calculate (33). We know that the dimension of all quantum registers involved is two, which means that $d_{O_{i}}=2M^{\prime}$. From (58), we get $\left\|\triangledown f_{\max}\right\|_{\infty}\leq a(p_{t})$. According to Condition 2 in Definition 3, for soundness, we require that, for every source $\phi$ and measurement device, Protocol 1 either certifies a minimum amount of GHZ entanglement or aborts with probability greater than $1-\varepsilon_{\operatorname{snd}}$ when applied on $\sigma$. So, the probability of Protocol 1 not aborting is negligible i.e., $\Pr[\Omega]_{\tau}\leq\varepsilon_{\operatorname{snd}}$. Also, note that $\varepsilon_{\operatorname{smo}}$ is chosen such that the fidelity constraint in (2) is satisfied. When Protocol 2 does not abort, we can define the following quantity, which corresponds to the right-hand side of (32):
$\eta(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},p_{t},\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},M^{\prime},M)\coloneqq\\ n\cdot f_{\max}\!\left(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},p_{t},M\right)+2\sqrt{n}\left(\log_{2}(1+2M^{\prime})+\lceil a(p_{t})\rceil\right)\\ \times\sqrt{1-2\log_{2}\left(\varepsilon_{\mathrm{smo}}\cdot\varepsilon_{\operatorname{snd}}\right)},$ (61)
@@
$\eta_{\mathrm{opt}}\left(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},M^{\prime},M\right)\coloneqq\\ \min_{p_{t}:p_{\min}<\frac{p_{t}(1)}{\gamma}<p_{\max}}\eta\left(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}},p_{t},\varepsilon_{\mathrm{smo}},\varepsilon_{\operatorname{snd}},M^{\prime},M\right).$ (62)
@@
For every source and all measurement devices in the setting detailed earlier, let $\rho$ be the state generated using Protocol 2, $\Omega$ the event that Protocol 2 does not abort, and $\rho_{|\Omega}$ the state conditioned on $\Omega$. Then, for all $8\cdot 3^{M/2}\sqrt{\varepsilon_{\operatorname{smo}}},\varepsilon_{\operatorname{snd}}\in(0,1)$, either Protocol 2 aborts with probability greater than $1-\varepsilon_{\operatorname{snd}}$ or
@@
Now, we have all the ingredients to show that Protocol 1 is sound.
@@ V.5 Soundness
In this subsection, we show that Protocol 1 is sound. To do this, we need to bound the quantity in (30) in a device-independent manner. We use the preceding theorems to obtain this bound and certify the multipartite entanglement distillation rate from a source. This brings us to our final theorem, which is as follows.
@@
Fix $\varepsilon_{\operatorname{smo}}>0$ such that $\varepsilon\coloneqq 8\cdot 3^{M/2}\sqrt{\varepsilon_{\operatorname{smo}}}\in(0,1)$. For all $\varepsilon_{\operatorname{snd}},\varepsilon_{\mathrm{cmp}},\varepsilon\in(0,1)$, Protocol 1 is a DIMEC protocol with
@@
    $E_{D}^{\varepsilon}(\rho_{|\Omega})\geq\\ \frac{-\eta_{\mathrm{opt}}\left(\omega_{\mathrm{exp}},\varepsilon_{\operatorname{snd}},\varepsilon_{\mathrm{smo}},M-1,M\right)}{M-1}+\frac{2\log\varepsilon_{\operatorname{smo}}}{M-1},$ (64)
@@
$E_{D}^{\varepsilon}(\rho_{|\Omega})\\ \geq\frac{-\eta_{\mathrm{opt}}\left(\omega_{\mathrm{exp}},\varepsilon_{\operatorname{snd}},\varepsilon_{\mathrm{smo}},M-1,M\right)}{M-1}+\frac{2\log\varepsilon_{\operatorname{smo}}}{M-1},$ (68)
@@
We plot also $-\eta_{\mathrm{opt}}(\omega_{\exp})/n(M-1)$ for $M=4$ and $\gamma=0.5$ up to leading order in Figure 1.
Figure 1: Plot of $-\eta_{\mathrm{opt}}(\omega_{\exp})/n(M-1)$, defined in (62), for $M=4$ and $\gamma=0.5$ up to leading order in $n$ for $\omega_{\exp}\in[p^{e}_{\min},p^{e}_{\max}]$ defined in (21). The solid line corresponds to $-\eta_{\mathrm{opt}}(\omega_{\exp})/n(M-1)$. The dashed line corresponds to $-\eta_{\mathrm{opt}}/n(M-1)=0$. The quantity $-\eta_{\mathrm{opt}}(\omega_{\exp})/n(M-1)$ represents the amount of multipartite distillable entanglement that can be certified per round by Protocol 1 as function of $\omega_{\exp}$.
@@ VI Conclusion
In this work, we defined a DI, $M$-partite entanglement distillation certification protocol consisting of completeness and soundness conditions. We then proposed a protocol and showed that it is indeed a DI multipartite entanglement distillation certification protocol. Specifically, we proved that Protocol 1 is indeed a DI multipartite entanglement distillation certification protocol. Our proposed protocol is centered around the MABK inequality [33, 34, 35], which is critical to proving the completeness and soundness conditions. To prove the soundness condition, we used the entropy accumulation theorem [36], the structure of the MABK inequality [33, 34, 35], and the multipartite entanglement distillation protocols and rates described in [41, 42].
For future works, we are interested in using other Bell-type inequalities that involve more than two binary measurements or measurements that have more than two outcomes. In this work, we have focused entirely on GHZ states and leaving out the W state. The W and GHZ states are not inter-convertible using LOCC alone [49]. Hence, our analysis here for GHZ states does not immediately apply to W states, but it is an interesting extension. It is also worth noting that there are infinitely many classes of genuinely multipartite entangled states that involve four or more parties [50]. It will be interesting to see what a unified certification protocol for such classes of states might look like.